OpenStudy (anonymous):

PLEASE HELP ME IM SO SERIOUS PLEASEEE SOMEBODY For the data in the table, does y vary directly with x? If it does, write an equation for the direct variation. (Hint: you need to find the “k” value) x 20 30 40 y 2 10 18 PLEASE HELP MEEEEEE

OpenStudy (anonymous):

so these are our points: (20, 2) (30, 10) (40, 18) make sense so far?

OpenStudy (anonymous):

YES @DemolisionWolf

OpenStudy (anonymous):

so for every 10 units in x, does the y change the same or is it a multiple?

OpenStudy (anonymous):

ITS MULTIPLE

OpenStudy (anonymous):

are you learning about exponential functions in you class right now?

OpenStudy (anonymous):

im not really sure my teacher just gave us the study guide.. non of these questions will be on the test but they will be similar so im trying to learn how to do it

OpenStudy (anonymous):

alright, give me a sec to see if i'm doing this right..

OpenStudy (anonymous):

okay ( :

OpenStudy (anonymous):

so I've looked it over, and the points given are linear, meaing you can solve this using the y=mx+b approach

OpenStudy (anonymous):

from 2 to 10, its +8 from 10 to 18, its +8 so for every 10 units in the x direction, y changes by 8 get it?

OpenStudy (anonymous):

yes i get it a little bit so how do i find the equation

OpenStudy (anonymous):

the equation will be in the form of y = mx + b we have x and y, so we need to find m and b, lets find m first, m is the slope, can you find the slope from one point to the next?

OpenStudy (anonymous):

is this right because in the hint it says find k how am im suppose to find k then do y = mx+b that sounds confusing

OpenStudy (anonymous):

in you class, what does k stand for? slope? y-intercept?

OpenStudy (anonymous):

this was out of the blue its a study guide for the exam so i have no clue

OpenStudy (anonymous):

i'm sorry, I believe I understand what the question is asking:

OpenStudy (anonymous):

its okay

OpenStudy (anonymous):

"does y vary directly with x?" answer: yes. "write an equation for the direct variation" answer: 10:8, because for every 10 units in the x direction, the y moves 8. (im not familiar with what k means, i'm sorry)